There are 2 girls and 7 boys in a math club. A team of four persons must be chosen for a tournament, and there must be at least 1 girl on the team. In how many ways can this be done?
Found 2 solutions by stanbon, sudhanshu_kmr:
Answer by stanbon(75887)

There are 2 girls and 7 boys in a math club. A team of four persons must be chosen for a tournament, and there must be at least 1 girl on the team. In how many ways can this be done?
-----
P(at least one girl on the team) = 1 - P(no girls on the team)
---
= 1 - 7C4/9C4
= 1 - 0.278
= 0.722
===============
Cheers,
Stan H.

Answer by sudhanshu_kmr(1152)

2 girls, 7 boys , total 9 persons
No. of ways to choose 4 persons from 9 = 9C4

no. of ways to choose 4 persons from 7 boys = 7C4 ( no girls)


according to question, there must be at least 1 girl, so we will reduce number of ways in which no girls in team...

total no. of ways in which there must be at least 1 girl = 9C4 -7C4
= 126 - 35

=91

if any problem related to the concept, u r welcome to contact ......